Two books cost rs150 altogether. One book costs rs30 more than the other. How much does each book cost?
Answer by Theo(13342)

let x = the price of the first book
let y = the price of the second book

x + y = 150

x = y + 30

replace x with y + 30 in the equation x + y = 150 to get:

y + 30 + y = 150

combine like terms to get 2y + 30 = 150

subtract 30 from both sides of the equa-tion to get 2y = 120

divide both sides of the equation by 2 to get y = 60

x = y + 30 becomes x = 90

one of the books costs 90.
the other book costs 60.